OpenStudy (anonymous):

What is the solution to the equation? 16+x-11+=18-5 x=_________ what does x equal

OpenStudy (anonymous):

@mathmath333 @Margaret123 @ScarlettGrace<3 @Here_to_Help15 @madison.bush

OpenStudy (stamp):

16 + x - 11 = 18 - 5 Add 11 to each side, what happens is: 16 + x = 18 - 5 + 11. Now subtract 16 from each side, what happens is: x = 18 - 5 + 11 - 16 Now you can find exactly what x is by doing 18 - 5 + 11 - 16.

OpenStudy (anonymous):

soo...the answer is 8 ?

OpenStudy (stamp):

corre ct

OpenStudy (anonymous):

can u help me with another one?